OpenStudy (anonymous):

What is the center and radius of the circle with equation (x + 9)2 + (y + 5)2 = 64?

hartnn (hartnn):

with, \(\Large (x-h)^2+ (y-k)^2 =r^2 \) the centre C = \(\large (h,k)\) and Radius = r

hartnn (hartnn):

does that help you in finding centre and radius of your question ?